find the dimensions of the rectangle field if it has a perimeter of 140 meter and an area if 1200 square meters.
Answer by Alan3354(69443)

find the dimensions of the rectangle field if it has a perimeter of 140 meter and an area if 1200 square meters.
--------------
P = 2L + 2W = 140
L + W = 70
L*W = 1200
---------
Find 2 factors of 1200 with a sum of 70.